what are magnitude and direction of electric field that will just support an electron so that it long in air


1
Expert's answer
2021-01-19T07:09:08-0500

There are two forces that act on the electron: the electric force "F_e" directed upward and the force of gravity "F_g" directed downward. In order to electron to be long in air, the force of gravity must be balanced by the electric force:


"F_e=F_g,""qE=m_eg,""E=\dfrac{m_eg}{q},""E=\dfrac{9.1\cdot10^{-31}\ kg\cdot 9.8\ \dfrac{m}{s^2}}{-1.6\cdot10^{-19}\ C}=-5.6\cdot10^{-11}\ \dfrac{N}{C}."


The magnitude of the electric field is "E=5.6\cdot10^{-11}\ \dfrac{N}{C}." The sign minus indicates that the direction of the electric field is downward.

Answer:

"E=5.6\cdot10^{-11}\ \dfrac{N}{C}," downward.